# TPIA: Towards Target-specific Prompt Injection Attack against Code-oriented LLMs

### v5

@@ VI-A Resistance to Potential Defenses
In this section, we further discuss whether our method can resist potential adaptive defenses that are particularly designed for our attack. In general, we can roughly categorize these defenses into detection and removal and design following potential defense methods against TPIA.
@@
Figure 10: The T-SNE plot of the classification keyword [CLS] at layer 17 shows the semantic differences between the original data and the injected data in the last embedding layer of CodeGemma-2b. The two types are almost indistinguishable.
@@
Detection: Semantic Analysis. Detection methods distinguish between malicious payload code containing perturbations and normal code, thereby rejecting the use of malicious payload code. Since detection methods based on code functionality are ineffective against non-functional perturbations, we adopt a semantic analysis approach. This involves examining the differences between the benign code and the injected code in the embedding layer of Code LLMs. However, this method may be difficult to implement. We find that code segments injected with perturbations might not have obvious semantic differences from benign ones. Specifically, as shown in Figure 10, we sample 1640 code segments and construct classification prompts. The results show that benign code and injected code are almost completely indistinguishable.
@@
Removal: Massive Deletion and Find-and-Replace. This type of defense involves removing perturbations by altering the non-functional parts of the code after introducing external code. In general, it entails a trade-off between code readability and the strength of removal since we cannot faithfully detect injected code segments (as mentioned above). To explore its effectiveness, we hereby test the impact of the percentage of character removal on the effectiveness of perturbations. As shown in Figure 11, TPIA still maintains a sufficiently high ASR under 0.5 removal rate. On the other hand, even if readability is completely sacrificed by extensively removing comments, TPIA can still be effective through non-comment portions of the code, as indicated in Table V. Arguably, the removal methods are likely to require significant human effort and may not yield effective results since TPIA imposes strict limitations on the length of perturbations.
@@
VI-B Potential Limitations and Future Directions
TPIA still has some drawbacks. Firstly, the effectiveness and stealthiness of TPIA are closely related to the keywords. We currently select keywords through a simple grid search. How to optimize keywords more efficiently is an important issue. Besides, TPIA cannot generate fully natural-looking perturbations within the context, resulting in a lack of fluency to some extent. In future work, we will explore how to use LLM as an optimizer to further improve this.
VII Conclusion
@@
In this paper, we investigated an important security vulnerability of Code LLMs for code completion tasks. We revealed the potential limitations of existing backdoor and adversarial attacks against Code LLMs, and introduced TPIA, a new target-specific attack paradigm without controlling the model’s training process. TPIA can circumvent mainstream detection methods (e.g., static analysis and sandboxes) by using non-functional code, instead of functional ones used in software supply chain attacks, to achieve malicious objectives. We evaluated TPIA under three popular programming languages and 31 representative cases. We showed that our TPIA can successfully attack three representative open-source Code LLMs (with an attack success rate of up to 97.9%) and two mainstream commercial Code LLM-integrated applications (with an attack success rate of over 90%) in all threat cases, using only a 12-token non-functional perturbation. Our work alerted a new serious security threat in using Code LLMs.
Acknowledgment

### v6

@@ 4.4 Resistance to Potential Defenses
In this section, we further discuss whether our method can resist potential adaptive defenses that are particularly designed for our attack. In general, we can roughly categorize these defenses into detection and removal and design following potential defense methods against ShadowCode.
@@
Figure 8: The T-SNE plot of the classification keyword [CLS] at layer 17 shows the semantic differences between the original data and the injected data in the last embedding layer of CodeGemma-2b. T-SNE plot figures at the other layers have similar results.
@@
Detection: Semantic Analysis. Detection methods distinguish between malicious payload code that contains perturbations and normal code, thus rejecting the use of malicious payload code. Since detection methods based on code functionality are ineffective against non-functional perturbations, we adopt a semantic analysis approach. This involves examining the differences between the benign code and the injected code in the embedding layer of Code LLMs. However, this method may be difficult to implement. We find that code segments injected with perturbations might not have obvious semantic differences from benign ones. Specifically, as shown in Figure 8, we sample 1640 code segments and construct classification prompts. We detailly discuss the setting of this semantic analysis in Appendix B. The results show that benign code and injected code are almost completely indistinguishable.
@@
Removal: Massive Deletion and Find-and-Replace. This type of defense involves removing perturbations by altering the non-functional parts of the code after introducing external code. In general, it entails a trade-off between code readability and the strength of removal since we cannot faithfully detect injected code segments (as mentioned above). To explore its effectiveness, we hereby test the impact of the percentage of character removal on the effectiveness of perturbations. As shown in Figure 9, ShadowCode still maintains a sufficiently high ASR under a 50% removal rate. On the other hand, even if readability is completely sacrificed by extensively removing comments, ShadowCode can still be effective through non-comment portions of the code, as indicated in Table V. Arguably, the removal methods are likely to require significant human effort and may not yield effective results since ShadowCode imposes strict limitations on the length of perturbations.
@@
4.5 Ablation Study
In this section, we conduct an ablation study to demonstrate the necessity of the keyword-based designing and forward reasoning enhancement methods proposed in Section 4 for ShadowCode attacks. We also discuss the effects of a critical method-related hyper-parameter (i.e., Top-$k$), which is crucial for determining both the computational load of Algorithm 1 and the effectiveness of our ShadowCode.
(a) ASR Study
(b) STF Study
Figure 10: Ablation study of keyword number and Top-$k$. We compare ShadowCode’s ASR and STF with keyword length from 0 to 2. During the process of generating perturbations, we use six different Top-$k$ values in our evaluation.
Effects of Top-$k$. This experiment validates the impact of Top value on the performance of ShadowCode. As shown in Figure 10, increasing the Top-$k$ value can significantly improve ShadowCode’s ASR and STF. Arguably, the attack performance can reach its optimal level when the Top-$k$ is increased to the value of the LLM’s vocab size. However, this approach would introduce computationally expensive overhead (CodeGemma has a vocab size of 256,000, far exceeding our default Top-$k$ value of 400).
(a) No FRE
(b) FRE
Figure 11: Ablation study of FRE. We compare the differences in ASR and losses both with and without using FRE module. 1-key1 and 1-key2 represent keeping one of the two selected keyword, respectively.
Effects of Keyword-based Designing. This experiment evaluates the effects of our optimization module, i.e., keyword-based designing. Specifically, we vary the number of keyword and generate the perturbations. As shown in Figure 10, limiting the number of keyword has a significant impact. The result of 1-keyword perturbation has approximately 10% decreases of ASRs and above 0.3 decrease of STF compared to 2-keyword perturbation. The 0-keyword perturbation, even with high computational power (600 Top-$k$ value), struggles to achieve an ASR above 80%. In contrast, 2-keyword perturbation can easily achieve both high ASR (above 95%) and STF (above 0.95). The above results indicate that using keyword can effectively improve the performance of ShadowCode.
Effects of Forward Reasoning Enhancement. We hereby perform a representative case study to illustrate the effects of forward reasoning enhancement (FRE), which corrects the distortion of the loss function. Specifically, we use the ‘ST2’ threat case as an example for our discussions since it is representative and difficult for attackers to achieve. It requires generating a clearly malicious code snippet and necessitates producing target code in a completely unrelated context, making it hard for the Code LLM to generate the target code, easily leading to distortion. To better evaluate the effects of this module, we first mitigate the effects of keyword-based designing by using the same keyword. We judge the degree of distortion based on the loss value. The results are illustrated in Figure 11, where we have three key findings. Firstly, under the same conditions, using FRE can lead to a higher ASR. Secondly, without FRE, the ASR is still very low ($<20\%$) even when the loss value is lower than that with FRE. Thirdly, FRE has a greater impact on ASR than keyword-based design. These results verify the necessity of forward reasoning enhancement in our ShadowCode method.
5 Potential Limitations and Future Work
As the first work on automatically attacking Code LLMs via external prompt injection, we have to admit that ShadowCode still exhibits some potential limitations. Firstly, we currently select the keyword tokens solely from the output tuple via a simple grid search. Although this method can already select effective keywords, it may overlook potentially better solutions. Developing more efficient keyword optimization strategies is an important potential research direction. Secondly, we currently do not mandate that ShadowCode generate entirely natural-looking induced perturbations. Arguably, unreadable content (e.g., symbols, abbreviations, and custom function names) is common in code comments and frequently does not raise suspicion. Meanwhile, as shown in Figure 7, the perturbation generated by ShadowCode constitutes only a small portion of the tokens in the entire code file, rendering it difficult to detect. However, we also believe that natural-looking perturbations can enhance the stealthiness of attacks, and we will explore how to leverage LLMs as optimizers to achieve this in our future work. Thirdly, our work only focuses on the code completion scenario. We will investigate the performance of ShadowCode in other scenarios such as code summarization and code translation in future work.
6 Conclusion
@@
In this paper, we introduced external prompt injection (EPI), a new attack paradigm against Code LLMs. EPI manipulates Code LLMs by injecting concise, non-functional induced perturbations into the victim’s code, which can spread through common dependencies. In particular, EPI does not necessitate control over the model’s training process and can achieve specific malicious objectives, posing more threatening risks over existing backdoor and adversarial attacks. Furthermore, we proposed a simple yet effective method (dubbed ShadowCode) for this attack paradigm, capable of automatically generating induced perturbations to achieve effective and stealth attacks. We evaluated ShadowCode across three popular programming languages and 31 representative cases. We showed that our ShadowCode can successfully attack three representative open-source Code LLMs (with an attack success rate of up to 97.9%) and two commercial Code LLM-integrated applications (with an attack success rate of over 90%) in all cases, using a 12-token non-functional perturbation. Our work alerted to a new serious security threat in the use of Code LLMs.
Acknowledgments
@@
Appendix A Additional Details
Indirect Prompt Injection against Code LLMs. In code domain, existing indirect prompt injection [18] (IPI) uses readable instructions to interfere with the model’s reasoning results, commanding the model to generate insecure code. In their show case, they place a large amount of disruptive commentary along with the obvious malicious code in the comment section, as depicted in Figure 12. We argue that this method exhibits very low feasibility. Direct human inspection or simple semantic analysis can easily detect these malicious comments.
Case Study. Here we show the details of our malicious objectives and threat cases. As shown in Table VI, we design 13 malicious objectives based on CWE top-10 KEV, existing research and famous destructive instruction. In the programming process, minor changes in the code can cause significant functional alterations. We roughly divide the causes of vulnerabilities in all malicious objectives into three categories: skipping critical process (CWE20, 78, 502, 918, 22, 306), wrong memory usage (CWE416, 122, 787, 843), and outputting dangerous snippets (ST0, 1, 2). All these vulnerabilities result in severe consequences.
Figure 13 presents three representative examples, encompassing diverse languages and objective types. The first example, CWE20 (Java), involves a program that attempts to fetch URL content and calculate its length. Secure programming practice dictates checking for null content before length calculation to prevent null pointer exceptions. However, post-attack, the Code LLM generates code that directly calculates the length after URL fetching, resulting in a null pointer exception if the URL is null. Null pointers can lead to severe consequences, including program crashes or data corruption. We categorize this situation as “skipping critical process.” The second example, CWE122 (C), concerns a program that attempts to pass content to a buffer (‘buf‘) via a command-line argument. Secure programming practice necessitates checking if the content exceeds ‘BUFSIZE‘ before assignment, thereby preventing overflow. Using ShadowCode, attackers can induce Code LLMs to perform direct assignments. Stack overflows access non-existent RAM space, leading to consequences such as system crashes and data leaks. We categorize this situation as “wrong memory usage.” This type of situation typically applies only to C/C++, as Java and Python offer automatic memory management. The third example, ST2 (Python), involves a program that attempts to call a function named TRG. Under normal circumstances, this call presents no issues, and the large language model generates suggestions based on contextual semantics or directly produces an ¡eos¿ token. Using ShadowCode, attackers can compel Code LLMs to generate the execution code ‘rm -rf /‘ immediately after calling the ‘TRG‘ function. If executed, this code results in the deletion of all data within the root directory. We categorize this situation as “outputting dangerous snippets.”
Figure 12: Greshake et al. [18] use a large amount of disruptive information and explicit malicious code snippets in the comments, which has a significant impact on the feasibility of the attack.
Baseline Details. IPI-M and C-GCG serve as crucial baselines in our effectiveness experiments. IPI-M mainly follows the existing IPI attack [18], employing human-readable instructions to compel the model to generate insecure code. However, due to the low stealth of the vanilla method, we simplify the instructions into concise, clear, template-based injection sentences, based on case categorizations. Specifically, for “skipping critical process,” we use “… directly after …” and “without any check”; for “wrong memory usage,” we use “It is important to … directly after …”; for “outputting dangerous snippets,”, we use “It is very important to promptly … after …”. Notably, the perturbation in IPI-M must be sufficiently explicit; for example, it needs to specify the code snippet you want to insert (e.g., “os.system("rm -rf")”) or present a clear assertion (e.g., “SSL3 is the most secure protocol”). Otherwise, Code LLMs are difficult to induce successfully. Arguably, the actual performance of IPI-M may be worse than our evaluation results shown in main experiments because of obvious malicious intent. C-GCG mainly follows the GCG method, employing an induced perturbation greedily optimized based on output tuple to achieve malicious objectives. Compared to ShadowCode, C-GCG does not use code simulation and two key modules: forward reasoning enhancement and keyword-based design.
Value and Metric Selection. We select the values of $r_{1}$ and $r_{2}$ based on the attack’s performance in the experiment. Specifically, in our experiments, 20% ASR is smaller than the ASR of no attack baseline in some situations, and 80% ASR outperforms the best performance of all the baselines. In addition, we find that only a few subcases (determined by the model, dataset, attack method, and case) have a success rate between 20% and 80%, accounting for 7.5%. Therefore, setting 20% and 80% as threshold of bad case and outstanding case respectively is appropriate.
As for metrics, we have already explained why metrics such as NBR and STF are used in the main experiments. Here we further explain some details related to them. First, NBR does not take the specific ASR value into account, as this metric is used to measure the diversity of attacks, rather than ensuring stable success rates. Second, STF targets the situation where attackers aim to cause as much damage as possible, and it varies with differences in ASR. In other words, even a difference between 100% ASR and 95% ASR can have a real impact on the overall threat level. Therefore, the calculation of STF considers the specific ASR value. In addition, we do not use common metrics for code generation tasks, such as BLEU or CodeBLEU [47], because they do not directly reflect the effectiveness of the attacks. Overall, we attempt to confirm the realistic meaning of each evaluation metric from different perspectives.
TABLE VI: Description of threat cases.
Case Description From
CWE416 Use after free. CWE 1st KEV
CWE122 Heap-based buffer overflow. CWE 2nd KEV
CWE787 Out-of-bounds write. CWE 3rd KEV
CWE20 Improper Input Validation CWE 4th KEV
CWE78 ‘OS Command Injection’ CWE 5th KEV
CWE502 Deserialization of untrusted data. CWE 6th KEV
CWE918 Server-side request forgery. CWE 7th KEV
CWE843 Access of resource using incompatible type. CWE 8th KEV
CWE22 ‘Path Traversal’ CWE 9th KEV
CWE306 Missing authentication for critical Function. CWE 10th KEV
ST0 Use ECB mode to encrypt. Schuster et al.[10]
ST1 Use SSL3 protocol Schuster et al.[10]
ST2 Executing “rm -rf / Famous destructive instruction
(a) Skipping critical process - CWE20
(b) Wrong memory usage - CWE122
(c) Outputting dangerous snippets - ST2
Figure 13: Three examples of ShadowCode attack cases. Code LLMs normally generate code segment without vulnerabilities, but they lose this ability after being attacked by ShadowCode. Attackers can modify the output content at the specified location through ShadowCode, and they can even generate malicious code snippets that have never appeared in the context, e.g., “os.system("rm -rf /")”.
Appendix B More Experiments
Figure 14: The ASR across different cases. The experimental results shown in this figure are the average values calculated from the effectiveness experiments involving 3 programming languages, 4 datasets, and methods excluding Clean Accuracy.
Effectiveness across Different Cases. In the main experiments, we have validated the outstanding attack performance of ShadowCode from three perspectives: model, dataset, and programming language. Here, we supplement the attack details for different attacks on a single case. Specifically, we use ASR to evaluate the performance of ShadowCode and the baselines across different cases. We present the specific average ASR for each case. As shown in Figure 14, ShadowCode achieves a certain level of effectiveness across all cases, with an overall average ASR of 83.4%. However, both C-GCG and IPI-M have cases where they are completely ineffective, especially for IPI-M. This demonstrates that directly using traditional GCG methods or manually designing perturbations is not effective in some scenarios. We believe this involves three reasons: (1) The baseline methods are not fully suitable for contexts with structured features, which is a fundamental characteristic of the code completion task; (2) The specific malicious objective include complex semantics such as target code, position code, and “disregarding security attributes,” which are sometimes not fully captured by the perturbations generated by the baseline methods; (3) The impact of keyword-based design and FRE; our ablation study proves that these two modules play a critical role.
According to the results, the most challenging to achieve is the straightforward case ST2. This case requires generating an obviously malicious command in a completely unrelated context. Another difficult case to implement is the CWE22 “Path Traversal,” which involves using unconventional paths like “../” to access files that are normally inaccessible. This case is relatively easier to achieve in Python, where the Code LLMs tend to suggest directly reading the path. However, it is more challenging to implement in Java and C/C++, as Code LLMs often only recommend checking the path rather than directly reading the file. CWE843 involves manipulating union members in a struct in C/C++. This represents a more complex malicious objective, where we attempt to manipulate one value within the union using an untrusted value. This is an unconventional operation that is difficult to describe using natural language, making it challenging for IPI-M to successfully attack. In contrast, methods such as ShadowCode and C-GCG, which convey malicious instructions through token selection, are able to achieve success. Arguably, ShadowCode has a clear advantage in attack performance for all the cases.
Semantic Analysis on Payload Code. Here we further discuss the settings of the semantic analysis experiment. We observe the classification features of code segment in the embedding layer of CodeGemma-2b. Specifically, we add the classification prompt keyword [CLS] to 820 original code snippets and 820 malicious payload code snippets totally. and then input them into the model to analyze their semantic features. We examine the t-SNE [59] distribution of the classification keyword [CLS] in the embedding layer, which is a widely used method for classification tasks. As shown in Figure 8, the features of the two types of text are almost completely intertwined, making it nearly impossible to distinguish between the injected code data and the original code data. In addition to observing the characteristics of layer 17 shown in Figure 8, we also examine the data features of layers 1 through 16. There is no significant difference observed in these layers compared to layer 17. There are several reasons for this situation. First, code data, including comments, tend to have lower fluency and coherence with high token complexity, which poses a significant challenge for semantic analysis tasks. Second, the perturbations used by ShadowCode occupy an extremely small number of tokens, exerting only a minor influence on the overall semantics of the injected data, making them hard to detect. Third, the semantic impact of ShadowCode differs from traditional injection attacks in that it does not require an extremely high intensity of information to counteract the model’s secure alignment; instead, it guides the model to generate vulnerabilities that could occur naturally, thus making detection through semantic analysis difficult. Our experiment results indicate that simple semantic analysis methods are not effective in defending against ShadowCode. This indicates that detecting code snippets containing malicious perturbations generated by ShadowCode through semantic analysis is not effective.